# Assessing survival models by interval testing

### v1

@@ 3.2 Example 2: Selecting a parametric survival model for challenging data
In this example, we fit seven standard parametric survival models. The survival data has a challenging shape, making it difficult for the models to provide good visual fit. Our aim is to use the interval test methods to determine whether any of the survival models are appropriate or whether flexible survival models are required.
Intervals selected based on censors
First, we perform the interval test with intervals chosen using censors as in Example 1. The 7 parametric models were fitted using flexsurv [15]. The final plots and tables were generated by the BITsurv package which is available here.
The test statistics for each parametric model are presented in Table 2. The PAVSI and TFT overall p-values are well above the 0.05 threshold for all of the survival models. This might lead us to believe that the models are providing good fit. However, as discussed in Appendix C, these test statistics are highly conservative when using the censor interval approach, meaning the overall p-values are not informative unless they indicate a rejection (less than 0.05.)
For each model, there are between 1 and 4 individual flags. However, these flags are primarily descriptive. Of greater interest are the Bonferroni rejections, of which there are none. This seems good, and from Table 2 alone we might believe that our curves provide good fit. However, this is an unfortunate result of using intervals based on censors, see poor fit in Figure 2. Due to the large number of censors, the majority of intervals are very small, which makes it difficult to detect poor fit within these intervals. For example, there are many small intervals for the last 5 months of the curve. It would be more informative to consider a single interval for those 5 months. Because of these limitations, a specified interval approach is required.
Figure 2: Part a: The KM curve (purple), censors (vertical dashed lines), and the fitted generalized gamma curve (black). Part b: The binomial interval test results for 10 evenly-spaced intervals. Part c: Numbers at risk.
@@
Exponential 0 3 0.468 0.854
Gamma 0 2 0.686 0.887
Generalised gamma 0 1 0.881 0.955
Gompertz 0 4 0.267 0.884
Log-logistic 0 1 0.881 0.977
Log-normal 0 2 0.696 0.968
Weibull 0 3 0.468 0.871
Table 2: Example 2: Interval test results using censor defined intervals. (Note that Bonferroni rejected intervals are also counted as individual flags.)
Specified intervals (10 evenly-spaced)
There are many ways these intervals could be chosen, and several approaches are briefly mentioned in the discussion. For the sake of brevity, we will outline one of the simpler approaches. The data is split into 10 evenly-spaced intervals (owing to the classical convention of the Hosmer-Lemeshow test.) These intervals start at time 0 and end at the largest censor time. That is, for $k=1,...,10$
@@
To simplify code implementation, p-values were calculated using the sinib R package [16], which uses a saddle point approximation for the distribution of the sum of binomials. The results are presented in Table 3, and the interval test plot for generalised gamma is presented in Figure 2.
@@
Table 3: Example 2: Interval test results using 10 evenly-spaced intervals.
As we can see, Table 3 tells a very different story to the previous interval selection approach. The overall PAVSI p-value is below 0.05 for all models except log-logistic and log-normal, indicating poor fit in 5 out of 7 models. Similarly, the overall TFT p-value rejects 5 out of 7 models and is below 0.07 in the final two models, which again raises concerns about poor fit. Finally, all of the parametric models have a Bonferroni rejection except log-normal, suggesting only the log-normal is a feasible model. We investigate the log-normal model further by presenting the results for each of its 10 intervals in Table 4.
@@
(0, 4.34] 18.76 11 0.0289 Accept Accept
(4.34, 8.67] 22.52 23 0.5534 Accept Accept
(8.67, 13.01] 17.90 26 0.9677 Accept Accept
(13.01, 17.34] 13.67 21 0.9710 Accept Accept
@@
(26.01, 30.35] 7.71 9 0.6947 Accept Accept
[…]
(34.69, 39.02] 5.02 5 0.5247 Accept Accept
(39.02, 43.36] 1.72 3 0.8279 Accept Accept
Table 4: Example 2: Interval test for the log-normal model with 10 evenly-spaced intervals.
From Table 4, we see that the interval resulting in an individual flag for the log-normal model was between times 30.35 and 34.69 with a p-value of 0.0087, which is above the 0.0025 threshold of a Bonferroni rejection. Furthermore, we note that the 4th interval had a p-value of 0.971. If this value was only 0.004 higher, then the log-normal would have received a second individual flag and thereby a PAVSI overall p-value rejection. Furthermore, the log-normal performed the worst of all 7 models for both AIC and BIC (not presented). For these reasons, as well as the overall TFT p-value of 0.0699 (close to 0.05) we can be justifiably suspicious of the log-normal model. However, we have no strong evidence to outright reject it.
In conclusion, due to the Bonferroni rejections, we would suggest that all presented models except for the log-normal are unfeasible. This leaves us with only the log-normal model remaining, which is somewhat restrictive and has also been flagged for concern as described above. For this reason, it is recommended that exploring flexible models such as splines and piecewise would be appropriate.
4 Simulation summary
A simulation experiment was performed to verify results under a true null hypothesis. Additional details are provided in Appendix C. The simulation experiment aimed to confirm our methodological assumptions with a simple but realistic example. For the approach where intervals were chosen using censoring, the test statistics had a very low type I error rate, suggesting that they are implausibly conservative in practice. For the 10 evenly-spaced interval approach, the test statistics performed as expected with a type I error rate close to 0.05. The Bonferroni test performed well in both approaches but was somewhat conservative for the censor approach. In conclusion, the two primary assumptions hold fairly well for the 10 evenly-spaced intervals approach. (The two primary assumptions are that censor intervals $I_{j}$ are uninformative of the event generating process and that the mid-point p-values are approximately uniform[0,1].) We also note that varying the maturity of the data had only a minor impact on the evenly-spaced interval approach.
5 Discussion
In terms of applying the interval test, two approaches were presented. In the first, intervals of interest are defined entirely by the censors. In the second, intervals of interest were specified as 10 evenly-spaced intervals up to the final censor. There are strengths and weaknesses to both approaches. We address this in the discussion, where the following topics are covered: the censor interval approach, how to specify intervals, the benefits of the 10 evenly-spaced interval approach, testing with limited patient numbers, and a conclusion.
In our first approach, intervals are selected by the data using the censor times. Unfortunately, this approach is generally unhelpful in describing regions with heavy censoring. This is because very small intervals are unlikely to be rejected due to low power. Additionally, with this approach, our test statistics are very conservative in standard cases (see Appendix C). On the other hand, we argue that in some cases (not presented) this approach is indicative of informative or unusual censoring. That is, it is very apparent if a lot of censoring has occurred, and furthermore, we can see if an unusual number of events occurred within these intervals. As a final point, this approach underlies the specified interval approach, so it will always provide helpful supporting evidence.
In the second approach, intervals can be specified. There are several ways this could be done. In our example, we provide one approach (10 evenly-spaced intervals). If submitting to a health technology assessment body, we recommend using this approach. This prevents intervals from being systematically chosen such that model selection favours a treatment’s efficacy. However, in terms of further research, there are certainly additional options for how the intervals are chosen. One option is automatic interval selection based on equal statistical power. Another topic of consideration is restricting chosen intervals to have boundaries that are censor times.
There are several benefits to the 10 evenly-spaced intervals approach compared to the censor interval approach. The 10 intervals are generally easier to interpret with plots. Furthermore, the approach is informative for areas with heavy censoring (such as the tail of the KM curve), which is often the area of greatest importance for long-term extrapolation. Additionally, overall p-values perform much better for this method. It is for these reasons that in general we recommend the fixed interval approach. However, we acknowledge that the censor interval approach is still valid and has its own benefits.
In situations with limited data, say under 20 patients, the tests can still be used. However, they will likely be under-powered and accept the fitted model. This is not necessarily a bad thing. When we look at data with only 20 patients, visual inspection of curve fit can be challenging, and after a subjective evaluation, we may think that flexible survival models are appropriate. That is, we have potentially over-fitted the data. The interval tests can prevent this by providing an objective evaluation. The tests will likely confirm that the observed data is feasible for simple models, indicating that flexible models are not required.
In conclusion, the interval tests provide additional information that can support the model selection process. The tests are able to highlight areas of particularly poor fit. It is recommended that they are especially useful in aiding decisions when the curve fit is poor.
6 Appendix
Appendix A In defense of midpoint p-values
When deriving test statistics and results, we will often make the assumption that midpoint p-values are uniform[0,1] under H0. If the validity of this approach is in question, then randomized p-values can be used, which are known to be uniform[0,1] under H0[13]. For this reason, the randomized p-values are superior from this purely statistical standpoint. However, we argue below that the midpoint p-values are more practical and that for our applications the uniform[0,1] assumption is often reasonable.
In the case of reducing the number of intervals to say 10, as is described in example 2, the generating process is the sum of binomials. If the number of patients is reasonable, one would expect these intervals to contain a reasonable number of feasible discrete outcomes. This means the uniform[0,1] assumption is more valid, compared to when there are only a small number of possible outcomes.
We also acknowledge that in practice, randomized p-values are more difficult to deal with. When an analysis is rerun, the p-values are rerandomized. This would then involve rerunning the analysis to determine the level of variation in the randomised p-values and their resulting test statistic. This is helpful. However, this approach leads you down the path of using the expected value of these reruns as your final quoted value. That is, the midpoint p-value, which is the expected value of randomized p-value.
Appendix B Extension to individual patient models
In the introduction, we mentioned that the most commonly used calibration methods rely on risk scores [4]. However, in our approach, risk scores were not considered. Furthermore, in our examples, all of our models inherently assume equal risk scores for each patient. It is, however, possible to extend our methodology to models where the risk scores are not assumed constant. This would involve amending Equation 1 such that each individual is considered separately for each interval. That is, every patient is assigned a specific probability of experiencing an event (as defined by the model) for that interval. Subsequently, Equation 1 would be replaced by a collection of $N_{I_{j}}$ Bernoulli random variables with different probabilities.
Appendix C Additional details on simulation
In practice, the parametric model never truly underlies the data, meaning type I errors don’t make logical sense. However, it does provide an interesting thought-experiment; if the observed data was generated by its fitted model, is the observed data behaving as you would expect? If the answer is yes, then the model is reasonable. If the answer is no, then the model is likely unreasonable.
The primary purpose of this investigation by simulation was to confirm that the methodology behaves as expected with respect to type I errors. The realistic model was described by an event generating process $T\sim exp(\lambda)$ and a censor generating process for $C$ given by
Finally, $Time=min(C,T)$; and if $T\leq C$, $Event=1$, else $Event=0$. Taking the time to be in months, we are saying some patients are heavily censored in the last 4 months of the trial, patients have an expected event time of $1/\lambda$, and some patients are censored throughout the trial.
For the censored interval approach, test statistics were shown to be unrealistically conservative. This is due to an inflated number of interval p-values that are close to 0.5, which occurs due to using the midpoint p-value approach. That is, most of the small intervals will yield a midpoint p-value of 0.5. These are then transformed for the TFT to give effective values of 1, which inflates the overall TFT p-value, pushing it closer to 1. In general, more censor intervals result in the tests being even more conservative. In our simulation, with 200 patients, $\lambda=1/{10}$, and 10,000 iterations the PAVSI type I error rate was 0.0041 and for TFT this was 0.0001. This is not ideal as it indicates that in practice these test statistics would be overly conservative. Thereby, both test statistics are not recommended in practice when using the censor interval approach.
For the 10 evenly-spaced intervals approach, the test statistics behaved as expected. When the number of patients increased, the test statistics appeared to converge to the expected values. Even with small numbers of patients, the test statistics remained reasonable. For example, with 50 patients, $\lambda=1/{10}$, and 10,000 iterations the PAVSI type I error rate was 0.0331 and for TFT this was 0.0146.
For the 10 evenly-spaced interval approach with 200 patients, the PAVSI type I error rate was 0.071 and for TFT this was 0.04. Note that here the PAVSI type I error rate is above 0.05. This is as expected. In particular, when assuming the interval p-values are uniform[0,1], the type I error is 0.08614. (This result can be shown by considering our approach, which uses 10 intervals, a 0.05 rejection level, and a test statistic obtained by using the mid-point p-value.) As the number of patients increases, our PAVSI test statistic appears to converge towards this value, 0.08614. That is, the midpoint p-values are behaving as uniform[0,1] random variables. Similarly, the TFT appears to converge towards 0.05 as the number of patients increases.
We make a quick digression here to discuss the PAVSI type I error converging towards 0.08614. If we were to select a different number of intervals, say 7, then the PAVSI type I error would converge towards 0.0444 instead. However, this is not ideal. With lower patient numbers, this value decreases even further due to the midpoint p-values. Thereby the 7 interval approach would always be conservative, which we argue is a bad property for model selection. The 10 evenly-spaced interval method approaches a type 1 error of 0.08614 as the patients increase. However, a type I error slightly above 0.05 isn’t necessarily a bad thing, especially in practice. Also, note that when using 10 intervals with a simulation of 100 patients, the PAVSI type I error rate was 0.0556, which is approximately what we want as it is close to 0.05.
The Bonferroni type I error rates were approximately 0.02 for the censored interval approach and 0.04 for the 10 evenly-spaced interval approach. These values appear to remain approximately stable as the number of patients changes. That is, approximately within $(0.03,0.05)$ for the 10 evenly-spaced interval approach and within (0.017,0.022) for the censor approach, when considering 50 to 2000 patients with $\lambda=1/{10}$.
Finally, we experimented with different levels of maturity in the data. In particular, we used $\lambda=1/{10}$ (mature data), $\lambda=1/{30}$ (somewhat mature), and $\lambda=1/{70}$ (immature). For these simulated data, the KM curves finished at approximate survival levels of 0.11, 0.48, and 0.73 respectively. In the censored interval approach, when the data was less mature, the test statistics performed worse. (They became even more conservative.) This is due to an increased amount of censoring in the less mature data. For 10 evenly-spaced interval approach, the different levels of maturity in the data had a somewhat minimal impact on the type I error rates. In particular, when weakening the data from mature to immature ($\lambda=1/{10}$ to $\lambda=1/{70}$), this would approximately result in a decrease of around 0.01 in the type 1 error rate for the PAVSI, TFT and Bonferroni tests.

### v2

@@ 3.2 Example 2: Selecting a parametric survival model for challenging data
In this example, the survival data has a challenging shape, making it difficult for models to provide good visual fit. We fit seven standard parametric survival models using flexsurv [15]. Our results and tables are then generated by the BITsurv package, which is available here. Our aim in Example 2 is to determine whether any of the seven survival models are appropriate, or whether survival models with additional flexibility are required. Both censor interval and specified interval approaches are used.
3.2.1 Intervals selected based on censors
First, we perform the interval test with intervals chosen using censors as in Example 1. The 7 parametric models were fitted, and the test statistics for each parametric model are presented in Table 2. The PAVSI and TFT overall p-values are well above the 0.05 threshold for all of the survival models. However, as will be discussed in our simulation section, these test statistics are often conservative for the censor interval approach.
For each model, there are between 1 and 4 individual flags. However, these flags are primarily descriptive. Of greater interest are the Bonferroni rejections, of which there are none. This seems good, and from Table 2 alone we might believe that our curves provide good fit. However, this is an unfortunate result of using intervals based on censors. Due to the large number of censors, the majority of intervals are very small, which makes it difficult to detect poor fit within these intervals. For example, there are many small intervals for the last 5 months of the curve. Instead, it would be more informative if we could consider a single interval for those final 5 months. To allow this, a specified interval approach is used.
@@ 3.2.1 Intervals selected based on censors
Exponential 0 3 0.456 0.820
Gamma 0 2 0.686 0.859
Generalised gamma 0 1 0.876 0.939
Gompertz 0 4 0.257 0.853
Log-logistic 0 1 0.876 0.968
Log-normal 0 2 0.686 0.954
Weibull 0 3 0.456 0.841
Table 2: Example 2: Interval test results using censor defined intervals.
3.2.2 Ten evenly spaced intervals (a specified interval approach)
There are many ways these intervals could be chosen, and several approaches are briefly mentioned in the discussion. Here, we will outline one of the simpler approaches. The data is split into 10 evenly spaced intervals (owing to the classical convention of the Hosmer-Lemeshow test). These intervals start at time 0 and end at the largest censor time. That is, for k=1,,10k=1,...,10
@@ 3.2.2 Ten evenly spaced intervals (a specified interval approach)
Applying this approach to the data yields the results presented in Table 3. Of the seven fitted curves, we present the generalised gamma results in Figure 2.
The results of Table 3 tell a very different story to the censor interval approach. First, the overall TFT p-value is below 0.05 for all models except log-normal, indicating poor fit in 6 out of 7 models. Similarly, the overall PAVSI p-value mostly agrees with the TFT, rejecting 5 out of 7 models. However, the PAVSI should be interpreted with caution when using the 10-fixed interval approach as it can inflate the type I error rate, see discussion Appendix B. Finally, all of the parametric models have a Bonferroni rejection except log-normal. The log-normal model is described in greater detail in Appendix C, where the results for each of its 10 intervals are given in Table 6.
In conclusion, due to the high number of Bonferroni and TFT rejections, we suggest that the majority of the presented models are unsuitable. In terms of the interval tests, the log-logistic and log-normal were the best performers. However, even these models had their limitations, with the log-normal model being the worst of all 7 models for both AIC and BIC (not presented). For these reasons, it is recommended that exploring flexible models such as splines and piecewise would be appropriate for this data.
@@
Table 3: Example 2: Interval test results using 10 evenly spaced intervals.
Figure 2: Example 2: Plot results for generalised gamma. Part a: The KM curve (purple), censors (vertical dashed lines), and the fitted generalised gamma curve (black). Part b: The binomial interval test results for 10 evenly spaced intervals. Part c: Numbers at risk.
4 Simulation
Due to the novelty of this approach, we will now pay particular attention to its properties under different simulated scenarios. Specifically, we focus on type I errors as this is the central concept about which the methodology is constructed.
4.1 Construction of simulation
We focus our attention on the exponential model as this allows us scope to explore other critical scenarios such as data maturity and patient numbers. Applications to additional scenarios (such as other parametric models) could be explored in future work.
Each simulation is described by the underlying event generation process: $T\sim exp(\lambda)$ and an underlying censor generating process for $C$ given by
The observable trial data is constructed as $Time=min(C,T)$; and $Event=1$ if $T\leq C$, else $Event=0$. This construction can be visualised as a 22 Month trial with heavy censoring in the last 4 months of the trial and infrequent censoring throughout.
In total, 12 data scenarios will be explored. For the number of patients in the fitted treatment arm, we consider 4 possible sizes, $N:50,100,200$ and $500$. For data maturity, we consider three levels: $\lambda=1/10$ (mature), $\lambda=1/30$ (moderate), and $\lambda=1/70$ (immature), where at the end of the trial period (22 months), these levels of maturity imply underlying survival rates of $0.11,0.48,$ and $0.73$ respectively.
For all 12 scenarios, 10,000 simulations were run for each interval method (10 evenly spaced and censor) and each p-value approach (randomised and midpoint), totaling 480,000 simulated trial datasets. The results are presented here, and a rerun confirmed that any written interpretations remained consistent.
4.2 Simulation results: randomised p-values
A core assumption of our methodology is that the intervals are uninformative of the event-generating process. Randomised p-values allow us to test this core assumption via simulation, without the added concern of midpoint p-values influencing the results.
For all simulation scenarios, both the 10 fixed-interval and censor-interval approaches behaved well with respect to type I error rates. That is, these rates were close to 0.05 for both the Bonferroni and TFT tests. This indicates that the core assumption holds well in all of our simulated scenarios for both the 10 fixed-interval and censor-interval approaches. Tables for these results are provided in Appendix D.
4.3 Simulation results: midpoint p-values
Unlike randomised p-values, the midpoint p-values are not uniform[0, 1] under H0. Instead, this is an assumption we must make for practical purposes, see discussion in Appendix A. Below, we investigate the impact this assumption has on type I errors for our simulation.
The censor interval approach did not perform as well as the 10 fixed interval approach. Results for the censor interval approach are presented in Table 4. The type I error rates for the Bonferroni tests were slightly conservative, being between 0.01 and 0.04 in most cases. However, for the censor interval approach, the TFT was very conservative, having a type I error of 0.0001 in all the cases where $\lambda=1/30$ or $\lambda=1/70$. We argue that this occurs because a large proportion of each simulated dataset’s intervals had an expected number of events close to 0, and the majority of the observations in these intervals were 0. This results in an artificially high number of p-values close to 0.5 when using the midpoint p-value approach, influencing the TFT and causing the conservative results. (The 10 fixed interval approach does not suffer from this limitation.)
The 10 fixed interval approach performed preferably to the censor approach, with results generally being more stable and less conservative. Results for the 10 fixed interval approach are presented in Table 5. Data maturity had a moderate impact on the results. However, patient numbers had a more distinct impact, with lower patient numbers generally resulting in more conservative tests (lower type I errors). As patient numbers increased, the type I error rates appeared to approach the targeted level of 0.05. Across the scenarios, Bonferroni remained fairly stable (type I error rates generally between 0.02 and 0.05), whereas the TFT was slightly more variable (generally between 0.01 and 0.05).
4.4 Simulation conclusion
We explored the type I error rates of our methodology under a variety of scenarios for an underlying exponential survival model. As discussed in Appendix A, randomised p-values are not recommended for practical applications, but simulating this approach helped to validate the assumptions of our methodology. When midpoint p-values are used, both the 10 fixed interval and censor interval approaches were inclined to be conservative for all scenarios of our simulation. The 10 fixed interval approach was generally more stable, with type I error rates being closer to the targeted rate of 0.05. Because of this, we would, in general, recommend 10 fixed intervals as the preferred approach. However, the censor interval approach can still be a useful supporting tool.
Table 4: Simulation type I error results: Censor interval approach, midpoint p-values
$\boldsymbol{\lambda=1/10}$ $\boldsymbol{\lambda=1/30}$ $\boldsymbol{\lambda=1/70}$
Number of patients Bonferroni TFT Bonferroni TFT Bonferroni TFT
50 0.0384 0.0253 0.0184 0.0001 0.0132 0.0001
100 0.0323 0.0091 0.0175 0.0001 0.0117 0.0001
200 0.0252 0.0009 0.0139 0.0001 0.0122 0.0001
500 0.0241 0.0001 0.0131 0.0001 0.0089 0.0001
For each of the 12 scenarios (data maturity and patient numbers), 10,000 datasets were simulated.
Table 5: Simulation type I error results: 10 fixed interval approach, midpoint p-values
$\boldsymbol{\lambda=1/10}$ $\boldsymbol{\lambda=1/30}$ $\boldsymbol{\lambda=1/70}$
Number of patients Bonferroni TFT Bonferroni TFT Bonferroni TFT
50 0.0325 0.0262 0.0219 0.0175 0.0245 0.0077
100 0.0390 0.0355 0.0336 0.0340 0.0221 0.0210
200 0.0436 0.0415 0.0421 0.0394 0.0336 0.0318
500 0.0491 0.0476 0.0440 0.0483 0.0410 0.0432
For each of the 12 scenarios (data maturity and patient numbers), 10,000 datasets were simulated.
5 Discussion
Two approaches for interval testing were presented. In our first approach, intervals were selected by the data using the censor times. Unfortunately, this approach is generally unhelpful in describing regions with heavy censoring. This is because very small intervals are unlikely to be rejected due to low power. Additionally, with this approach, our simulations indicated that the test statistics are conservative in standard cases. However, we argue that the censor interval plot is still a useful tool. Specifically, by visually inspecting the censor interval plot and test results, one is more supported when informally deducing complex properties of the data.
In the second approach, we extended the methodology so that intervals can be specified. Specifically, we focused on the 10 evenly spaced interval approach. The 10 intervals were generally easier to interpret with plots compared to the censor interval approach. Additionally, the approach is informative for areas with heavy censoring, such as the tail of the KM curve, which is often the area of greatest importance for long-term extrapolation. Furthermore, our simulations indicated that this method generally performed better than the censor interval approach, especially if the number of patients exceeded 100. (That is, the type I errors were closer to the targeted 0.05 level.) For these reasons, we would generally recommend the 10 evenly spaced interval approach to the censor interval approach, especially for cases where the approaches are being used to formally accept or reject fitted survival models. To facilitate the application of these approaches, a package has been developed.
Regarding further research, there are several avenues for further exploration. Although we focused on 10 evenly spaced intervals, the methodology is flexible to accommodate many methods of interval selection. One possible extension would be automatic interval selection based on equal statistical power. Another extension could involve restricting chosen intervals to have boundaries that are censor times. Another avenue of exploration is to apply our methodology to risk score models. That is, more complex survival models that allow unequal risk scores for each patient (our examples inherently assumed equal risk scores). This approach would require some modifications to the methodology, specifically using a Poisson binomial distribution rather than the binomial distribution (Equation 1).
Our approach provides several advantages over standard model selection techniques such as AIC/BIC. First, the TFT and Bonferroni provide objective tests, which can be used to reject a fitted survival model. Moreover, the methods indicate areas of particularly poor fit (for example, any one of 10 evenly spaced intervals up to the end of the KM curve). Such information is invaluable to analysts, helping them to decide which survival models are most suitable.
Appendix A Midpoint p-values discussion
When deriving test statistics and results, we will often make the assumption that midpoint p-values are uniform[0,1] under H0. If the validity of this approach is in question, then randomized p-values can be used, which are known to be uniform[0,1] under H0 [13]. For this reason, the randomized p-values are superior from this purely statistical standpoint. However, we argue below that the midpoint p-values are more practical and that, for our applications, the uniform[0,1] assumption is often reasonable.
In the case of reducing the number of intervals to 10, as is described in example 2, the generating process is the sum of binomials. If the number of patients is reasonable, one would expect these intervals to contain a reasonable number of feasible discrete outcomes. This means the uniform[0,1] assumption is more valid compared to when there are only a small number of possible outcomes.
We also acknowledge that in practice, randomized p-values are more difficult to deal with. When an analysis is rerun, the p-values are rerandomised. This would then involve rerunning the analysis to determine the level of variation in the randomised p-values and their resulting test statistic. This is helpful. However, this approach leads you down the path of using the expected value of these reruns as your final quoted value. That is, the midpoint p-value, which is the expected value of the randomised p-value.
Appendix B PAVSI - practical discussion
In this paper, we focused primarily on the TFT test statistic as it utilises the full range of the data. That is, for PAVSI, a p-value of say 0.026 would contribute equivalently to a p-value of 0.5, which is arguably a limitation of this approach compared to the TFT. For this reason, in general, we recommend the TFT over the PAVSI.
Additionally, there is a caveat to using the PAVSI for the 10 evenly spaced interval approach when using midpoint p-values. Specifically, even if the p-values are uniform[0,1] under H0, when we use the midpoint p-value to obtain the PAVSI of the 10 evenly spaced intervals, it can be shown analytically that the type I error is 0.08614 rather than the targeted 0.05 level. Because of this limitation, the PAVSI results are not described in the simulation section. Regardless, interested readers can find these simulation results on Github. Additionally, PAVSI is included as an option in the available codes.
In conclusion, for the censor interval approach, the PAVSI behaves well and can be well used. For the 10 evenly spaced interval approach, the PAVSI should be used with caution as the type I error rate can be inflated. In real-world analyses, extreme observations in an interval (that is likely very small) can render the TFT useless on account of a single extreme p-value, which results in the TFT rejecting almost every model. For such cases, PAVSI is the recommended alternative. For standard cases, the TFT is generally preferred.
Appendix C Additional details for Example 2: Log normal model with 10 evenly spaced intervals
In Example 2, we looked at seven fitted survival models using a 10 evenly spaced interval approach. Of those seven models, only the log-normal model was not rejected by the Bonferroni test. Because of this, we explore the log-normal model in further detail. Specifically, we now look at the results in each of the 10 intervals for the log-normal model, Table 6.
From Table 6, we see that the only interval resulting in an individual flag for the log-normal model was between times 30.35 and 34.69 with a p-value of 0.0061, which is above the 0.0025 threshold of a Bonferroni rejection. Furthermore, we note that the 4th interval had a p-value of 0.9727. If this value was only 0.0023 higher, then the log-normal would have received a second individual flag, and as a result of this the PAVSI overall p-value would have yielded a rejection (noting caveat Appendix B). Furthermore, the log-normal performed the worst of all 7 models for both AIC and BIC (not presented). For these reasons, as well as the overall TFT p-value of 0.0513 (very close to 0.05) we can be justifiably suspicious of the log-normal model. However, we have no strong evidence to outright reject it.
Table 6: Example 2: Interval test details for the log-normal model with 10 evenly spaced intervals.
@@ Appendix C Additional details for Example 2: Log normal model with 10 evenly spaced intervals
(0, 4.34] 18.67 11 0.0301 Accept Accept
(4.34, 8.67] 22.55 23 0.5514 Accept Accept
(8.67, 13.01] 17.83 26 0.9690 Accept Accept
(13.01, 17.34] 13.59 21 0.9727 Accept Accept
@@
(26.01, 30.35] 7.69 9 0.6975 Accept Accept
(30.35, 34.69] 6.51 1 0.0061 Flag Accept
(34.69, 39.02] 4.95 5 0.5364 Accept Accept
(39.02, 43.36] 1.56 3 0.8598 Accept Accept
Appendix D Simulation results when using randomised p-values
As mentioned previously, the type I error rates for both Tables 7 and 8 were close to 0.05 for both the Bonferroni and TFT tests. However, it is noted that for the censor interval approach with $\lambda=1/10$ and $N=50$ the TFT and Bonferroni rates were slightly higher being close to 0.07, Table 7. This may indicate that for small N with a high rate of events our assumption (that the intervals are uninformative of the event‐generating process) may not hold as well compared to other cases.
Table 7: Simulation type I error results: Censor interval approach, randomised p-values
$\boldsymbol{\lambda=1/10}$ $\boldsymbol{\lambda=1/30}$ $\boldsymbol{\lambda=1/70}$
Number of patients Bonferroni TFT Bonferroni TFT Bonferroni TFT
50 0.0692 0.0713 0.0557 0.0584 0.0500 0.0524
100 0.0535 0.0635 0.0513 0.0541 0.0465 0.0562
200 0.0547 0.0601 0.0521 0.0540 0.0521 0.0526
500 0.0531 0.0557 0.0493 0.0511 0.0494 0.0499
For each of the 12 scenarios (data maturity and patient numbers), 10,000 datasets were simulated.
Table 8: Simulation type I error results: 10 fixed interval approach, randomised p-values
$\boldsymbol{\lambda=1/10}$ $\boldsymbol{\lambda=1/30}$ $\boldsymbol{\lambda=1/70}$
Number of patients Bonferroni TFT Bonferroni TFT Bonferroni TFT
50 0.0510 0.0550 0.0528 0.0561 0.0527 0.0549
100 0.0537 0.0518 0.0502 0.0504 0.0523 0.0554
200 0.0494 0.0543 0.0512 0.0556 0.0470 0.0539
500 0.0512 0.0492 0.0474 0.0487 0.0456 0.0532
For each of the 12 scenarios (data maturity and patient numbers), 10,000 datasets were simulated.